# HumanGif: Single-View Human Diffusion with Generative Prior

### v2

@@ Abstract
Previous 3D human creation methods have made significant progress in synthesizing view-consistent and temporally aligned results from sparse-view images or monocular videos. However, it remains challenging to produce perpetually realistic, view-consistent, and temporally coherent human avatars from a single image, as limited information is available in the single-view input setting. Motivated by the success of 2D character animation, we propose HumanGif, a single-view human diffusion model with generative prior. Specifically, we formulate the single-view-based 3D human novel view and pose synthesis as a single-view-conditioned human diffusion process, utilizing generative priors from foundational diffusion models to complement the missing information. To ensure fine-grained and consistent novel view and pose synthesis, we introduce a Human NeRF module in HumanGif to learn spatially aligned features from the input image, implicitly capturing the relative camera and human pose transformation. Furthermore, we introduce an image-level loss during optimization to bridge the gap between latent and image spaces in diffusion models. Extensive experiments on RenderPeople and DNA-Rendering datasets demonstrate that HumanGif achieves the best perceptual performance, with better generalizability for novel view and pose synthesis. Our code will be available at https://github.com/skhu101/HumanGif.
Figure 1: HumanGif is a single-view human diffusion model that leverages generative priors to compensate for information missing from the input image. With generative prior, HumanGif synthesizes perpetually realistic, view-consistent, and temporally coherent human avatars.
@@ 1 Introduction
Synthesizing 3D human performers with consistent novel views and poses holds extensive utility across various domains, including AR/VR, video games, and movie production. Recent methods enable the novel view and pose synthesis of 3D human avatars from sparse-view human videos [88, 19, 90, 125, 85, 119, 109, 59, 61, 114, 50, 21, 45, 48, 46], with Neural Radiance Field [82] or Gaussian Splatting-based [66] representation. While impressive novel view and pose synthesis results can be achieved from sparse-view videos, generating perpetually realistic, view-consistent, and temporally coherent human avatars from a single image [49, 29, 120, 75, 53, 86, 116, 42, 20, 15, 14] remains challenging due to the limited information available in this setting.
@@
To address the challenge of complementing information missing from a single image, a line of research [120, 53, 86, 14] focuses on the novel view synthesis from a single human image by complementing human images or UV textures at unseen views (e.g., back view) through additional models (e.g., T2I diffusion models). To further learn view-consistent and temporally aligned human avatars from a single image, another line of research [49, 105] proposes synthesizing novel views and poses from a single image through learning a generalizable HumanNeRF or a human diffusion model from scratch. However, such frameworks normally require multi-view human image collections as training datasets, and their performance is closely tied to the quality and scale of these datasets. While efforts to create high-quality multi-view human image datasets [131, 38, 22, 73, 11, 89, 41, 110, 4, 56, 80, 90, 36, 37, 55, 57, 128] have accelerated in recent years, the number of human subjects in these datasets remains significantly smaller compared to multi-view object and scene image datasets [27, 122, 133, 111, 26, 142]. This data disparity limits the generalization performance of single-image-based 3D human modeling frameworks.
@@
To mitigate this data-sparsity dilemma, recent research [9, 65, 115, 18, 126, 47, 143, 137, 67, 76] in 2D character animation involves generative prior (inherit pre-trained weights) from Text-to-image (T2I) diffusion models (e.g., Stable Diffusion [96]) to assist the single-view based 2D human animation (novel pose synthesis). By inheriting the generative priors from diffusion models, these approaches demonstrate impressive generalization performance for novel human subjects and poses in 2D character animation tasks (e.g., dancing video generation) using only a modest amount of training data, such as hundreds of monocular dancing videos [58]. Motivated by these successes, it is desirable to explore the use of generative prior from T2I diffusion models in the single-view-based 3D human novel view/pose synthesis task.
@@
In this work, we propose HumanGif, a single-view conditioned human diffusion with generative prior. Specifically, we formulate the single-view-based 3D human novel view and pose synthesis as a single-view-conditioned human diffusion process, utilizing generative priors from foundational diffusion models to complement the missing information. Yet, this promising avenue comes with two primary barriers: 1) How to integrate the relative camera and human pose information into the human diffusion framework? 2) How to ensure that the performance achieved in the latent space is equally effective in the image-level space for human diffusion models? In particular, one potential solution is to allow diffusion models to learn relative camera and human pose transformation information through cross-attention between the reference image and the target pose image. However, our experiments reveal that human diffusion models struggle with a) learning detailed information (e.g., logos on T-shirts) even when such details are present in the reference image, and b) achieving view-consistent and temporally aligned results. To tackle the challenges, we draw inspiration from prior 3D reconstruction methods [49, 121] and introduce a Human NeRF module [49] to capture the explicit fine-grained features in the target space by transforming the human subject from the reference pose space to the target pose space using linear blend skinning (LBS) of SMPL [77] and integrating them with volume rendering. With the SMPL LBS transformation, the resulting rendered feature map provides a spatially aligned conditioning signal that encodes the relative camera and human pose transformation. To further enhance the view and temporal consistency, we incorporate the multi-view attention and temporal attention in our HumanGif. Additionally, we observe a discrepancy between the latent space and image space for human images, attributed to the Variational Autoencoder (VAE) [69] used in diffusion models. Inspired by the Radiance field rendering loss employed in 3D object diffusion models [83], we construct an image-level loss by mapping the noise latent to the image space through a VAE decoder. This ensures consistent optimizations in both latent and image spaces. With the above modules incorporated, our HumanGif successfully recovers fine-grained information from the reference image and achieves the best perceptually realistic, view-consistent, and temporally coherent results. Our main contributions are as follows:
@@
    To learn perpetually realistic, view-consistent, and temporally aligned 3D human avatars, we integrate a Human NeRF module with multi-view attention and temporal attention.
@@
    Extensive experiments on two human datasets demonstrate that HumanGif outperforms baseline methods in perceptual quality of novel view and pose synthesis.
@@ 2 Related Work
Novel View/Pose Synthesis from a Single Human Image. Although 2D human modelling [2, 71, 104, 81, 31, 62, 63, 32] has made substantial progress, it is still challenging to synthesizing 3D humans from monocular inputs, especially for a single human image input [100, 101, 39, 74, 30, 72, 10, 127, 7, 8, 54, 40, 141, 124, 123, 5, 25]. To complement the information missing from a single image input, these approaches [120, 53, 86, 116, 20, 23, 138, 42, 29, 15, 14] typically leverage text-to-image (T2I) diffusion models, such as Stable Diffusion [96], by employing subject-specific Score Distillation Sampling (SDS) [91] or fine-tuning a T2I diffusion model conditioned on a front-view image using multi-view human datasets or 3D scan datasets. This enables the synthesis of human images from unseen viewpoints, such as generating a back view from a given front view. To further learn view-consistent and temporally aligned human avatars from a single image, recent research [49, 105] leverages the strong modeling capabilities of deep neural networks to learn generalizable features for Neural Radiance Field (NeRF) or diffusion models, enabling the synthesis of novel views and poses in a feed-forward manner. However, the scale of multi-view human datasets limits their generalization ability in novel view and pose synthesis. In this work, we propose to utilize the generative prior from diffusion models to produce perpetually realistic, view-consistent, and temporally coherent human avatars from a single image.
@@
Diffusion Models for 2D Character Animation. 2D Character Animation aims at generating temporally coherent animation videos from one or more static human images [3, 12, 16, 31, 60, 92, 95, 103, 106, 107, 130, 132, 135, 140]. With advancements in text-to-image diffusion models, recent studies [9, 65, 115, 18, 126, 47, 143, 137, 67, 76] have investigated the use of diffusion models for 2D character animation. These approaches normally use a reference human image and a sequence of target pose images (e.g., OpenPose [13], DWPose [129], DensePose [34], or SMPL pose [77]) as conditional inputs to pose encoders or ControlNet [134], generating a sequence of target images aligned with the reference human image and target pose images. Motivated by the success of diffusion-based methods in 2D character animation, we adapt these methodologies for 3D/4D human modeling from a single image. Human4DiT [105] introduced a diffusion model that generates multi-view human animation videos from a single image, utilizing factorized image, view, and temporal modules, similar to our approach. However, the code for this model is not publicly available. We identified that our baseline multi-view video diffusion model could not capture fine-grained details and maintain multi-view consistency. We show that integrating generative prior and our Human NeRF module into the diffusion pipeline effectively resolves these challenges. Our concurrent work GAS [78] also utilizes diffusion models to enhance the novel view and pose synthesis performance from a Human NeRF module.
@@ 3 Our Approach
Our proposed HumanGif, as illustrated in Fig. 2, learns a single-view conditioned human diffusion model for novel view and pose synthesis by utilizing generative prior from Stable Diffusion. Specifically, given a reference human image and a target human pose sequence (estimated from a human video or generated from other modalities like text, or audio), HumanGif aims to predict a sequence of human pose images aligned with the person in the reference image and motions observed in the target human pose sequence. We assume the camera parameters are pre-calibrated and the human region masks can be estimated. We also assume the corresponding SMPL pose and shape parameters $\bm{\theta},\bm{\beta}$ are given.
Figure 2: HumanGif Framework. Given a single input human image and a target pose sequence, our HumanGif produces a sequence of target images aligned with the human subject in the input image and target human poses. To synthesize view-consistent and pose-consistent outputs, our HumanGif proposes to incorporate generative prior, a Human NeRF module, image-level loss, multi-view attention, and temporal attention.
@@ 3.1 Preliminary
3.2 Single-view conditioned Human Diffusion Model
@@
Pose Encoder. There are various options for defining target poses, including OpenPose, DensePose, DWPose, and SMPL parametric poses. We follow [143] to estimate SMPL mesh for the reference image, animate the SMPL mesh with the target SMPL pose parameters, and render normal images as the target guidance signal. The SMPL normal pose images are encoded through a Pose Encoder [47], which contains four convolution layers. The Pose Encoder output is added to the noise latent to provide view and pose guidance information.
Table 1: Quantitative comparison of our HumanGif and baseline methods on the RendePeople and DNA-Rendering datasets.
@@
Human NeRF. The proposed framework aims to generate consistent novel views and poses from a single human image. However, we observe two challenges in the learned human diffusion models: (a) difficulty in capturing fine-grained details (e.g., logos on T-shirts), even when they are present in the reference image, and (b) limited capability to maintain consistency across multiple views and poses. One underlying reason for the challenges is that the reference image is first encoded into a latent representation and then processed by ReferenceNet to extract features, resulting in information loss during this process. To tackle this challenge, inspired by prior research in Human NeRF [89, 49, 19, 87, 125, 85, 118, 109, 59, 61, 114, 70, 33, 23, 139, 52], we integrate a Human NeRF module into human diffusion models. This module learns fine-grained features that are spatially aligned with both the reference and target pose space, enhancing detail preservation and consistency.
@@
As illustrated in Fig. 2, the input to the Human NeRF module is a single human image $\mathbf{I}^{\text{ref}}$ along with its corresponding camera parameters $\bm{P}^{\text{ref}}$ and SMPL pose parameter $\bm{\theta}^{\text{ref}}$ and shape parameter $\bm{\beta}^{\text{ref}}$. The module outputs the rendered human feature image in the target camera view $\bm{P}^{\text{tgt}}$, corresponding to the target SMPL pose $\bm{\theta}^{tgt}$ and shape $\bm{\beta}^{\text{tgt}}$. Specifically, in the target space, we cast rays passing through the camera origin and image pixels, and sample points $\bm{x}^{\text{tgt}}$ along the cast rays. These points $\bm{x}^{\text{tgt}}$ are transformed into the canonical space $\bm{x}^{c}$ using inverse Linear Blend Skinning (LBS). Subsequently, hierarchical 3D-aware features,i.e., point-level features and pixel-aligned features, are queried from their respective feature extraction modules. The queried features are concatenated and passed to the NeRF decoder to predict the density $\sigma$ and feature $\bm{c}$ for each sampled point. The final pixel features are rendered in the target space through volume rendering, integrating the density and feature values of the sampled 3D points along the rays in the target space.
The details of extracting point-level features and pixel-aligned features are described as follows. 1) Point-level Features. We first extract per-point features by projecting the SMPL vertices onto the 2D feature map of the input image, as illustrated in Fig. 2. Next, we apply inverse LBS to transform the posed vertex features into the canonical space. These transformed features are then voxelized into sparse 3D volume tensors and further processed using sparse 3D convolutions [24]. From the encoded sparse 3D volume tensors, we extract point-level features $\bm{f}_{\text{point}}(\bm{x}^{c})$ for each point $\bm{x}^{c}$. With the awareness of 3D Human structure, point-level features capture local texture details in the seen area and infer textural information for unseen areas through sparse convolution. 2) Pixel-aligned Features. Due to limited SMPL mesh and voxel resolution, point-level features suffer from significant information loss, especially in areas visible from the reference image. To compensate for the information loss problem, we additionally extract pixel-aligned features by projecting 3D deformed points $\bm{x}^{c}$ into the input view. As illustrated in Fig. 2, each deformed point $\bm{x}^{c}$ is transformed into the observation space as $\bm{x}^{\text{ref}}=\text{LBS}(\bm{x}^{c};\bm{\theta}^{\text{ref}},\bm{\beta}^{\text{ref}})$ using LBS. It is then projected onto the input view, allowing us to query the pixel-aligned features. Leveraging the complementary strengths of point-level and pixel-aligned features, our Human NeRF module effectively captures fine-grained feature details in regions visible in the reference view while inferring features for regions occluded in the reference view.
@@
Multi-view/Temporal Module. To learn view-consistent and temporal coherent 3D human avatars, we further integrate a view-attention and temporal attention layer after the spatial-attention and cross-attention components within the Res-Trans block of Denosing UNet. Inspired by the efficient temporal attention layer adopted in [47, 35], we utilize the same architecture for our view-attention and temporal attention layer. The view-attention layer performs the attention in the view channel while the temporal attention layer performs the attention in the human pose channel.

### v3

@@ Abstract
Previous 3D human creation methods have made significant progress in synthesizing view-consistent and temporally aligned results from sparse-view images or monocular videos. However, it remains challenging to produce perpetually realistic, view-consistent, and temporally coherent human avatars from a single image, as limited information is available in the single-view input setting. Motivated by the success of 2D character animation, we propose HumanGif, a single-view human diffusion model with generative prior. Specifically, we formulate the single-view-based 3D human novel view and pose synthesis as a single-view-conditioned human diffusion process, utilizing generative priors from foundational diffusion models to complement the missing information. To ensure fine-grained and consistent novel view and pose synthesis, we introduce a Human NeRF module in HumanGif to learn spatially aligned features from the input image, implicitly capturing the relative camera and human pose transformation. Furthermore, we introduce an image-level loss during optimization to bridge the gap between latent and image spaces in diffusion models. Extensive experiments on RenderPeople, DNA-Rendering, THuman 2.1, and TikTok datasets demonstrate that HumanGif achieves the best perceptual performance, with better generalizability for novel view and pose synthesis. Our code is available at https://github.com/skhu101/HumanGif.
Figure 1: HumanGif is a single-view human diffusion model. By inheriting generative prior, HumanGif synthesizes realistic view and pose-consistent images.
@@ 1 Introduction
Synthesizing 3D human performers with consistent novel views and poses holds extensive utility across various domains, including AR/VR, video games, and movie production. Recent methods enable the novel view and pose synthesis of 3D human avatars from sparse-view human videos [93, 19, 95, 131, 90, 125, 115, 61, 63, 120, 52, 21, 47, 50, 48, 76], with Neural Radiance Field [87] or Gaussian Splatting-based [68] representation. While impressive novel view and pose synthesis results are achieved from sparse-view videos, generating perpetually realistic, view-consistent, and temporally coherent human avatars from a single image [51, 29, 126, 80, 55, 91, 122, 44, 20, 15, 14] remains challenging as limited information is available.
@@
To address the challenge of complementing information missing from a single image, a line of research [126, 55, 91, 14] focuses on the novel view synthesis from a single human image by complementing human images or UV textures at unseen views (e.g., back view) through foundational models (e.g., T2I diffusion models). To further learn view-consistent and temporally aligned human avatars from a single image, another line of research [51, 110] proposes synthesizing novel views and poses from a single image by learning a generalizable HumanNeRF or a human diffusion model from scratch. However, such frameworks normally require multi-view human image collections as training datasets, and their performance is closely tied to the quality and scale of these datasets. Although efforts to create high-quality multi-view human image datasets [137, 39, 22, 78, 11, 94, 43, 116, 4, 58, 85, 95, 37, 38, 57, 59, 134] have been accelerated in recent years, the number of human subjects in these datasets remains significantly smaller in comparison with the size of multi-view object and scene image datasets [27, 128, 139, 117, 26, 148]. This data disparity limits the generalization performance of single-image-based 3D human modeling frameworks.
@@
To mitigate the data-sparsity dilemma, recent research [9, 67, 121, 18, 132, 49, 149, 143, 69, 81] in 2D character animation involves generative prior (inherit pre-trained weights) from Text-to-image (T2I) diffusion models (e.g., Stable Diffusion [101]) to assist the single-view based 2D human animation (novel pose synthesis). By inheriting the generative priors from diffusion models, these approaches demonstrate impressive generalization performance for novel human subjects and poses in 2D character animation tasks (e.g., dancing video generation) using only a modest amount of training data, such as hundreds of monocular dancing videos [60]. Motivated by these successes, it is desirable to explore the use of generative prior from T2I diffusion models in the single-view-based 3D human novel view/pose synthesis task.
@@
In this work, we propose HumanGif, a single-view conditioned human diffusion with generative prior. Specifically, we formulate the single-view-based 3D human novel view and pose synthesis as a single-view-conditioned human diffusion process, utilizing generative priors from foundational diffusion models to complement the missing information. Yet, this promising avenue comes with two primary barriers: 1) How to spatially and temporally align the learned human avatar with the reference image and target human poses? 2) How to ensure the performance achieved in the latent diffusion space is equally effective in the image-level space? In particular, one potential solution to the first challenge is to inject cross-attention modules into diffusion models to learn spatial and temporal transformation from the reference image and target pose images. However, our experiments reveal that human diffusion models built for 2D character animation struggle with a) learning detailed information (e.g., logos on T-shirts) even when such details are present in the reference image, and b) achieving view-consistent and temporally aligned results.
To address the first challenge, we introduce a more spatially and temporally aligned conditioning signal drawing inspiration from 3D human reconstruction methods [125, 51, 127]. Specifically, we learn a Human NeRF module to transform the human subject from the reference pose space to the target pose space using a parametric human SMPL model [82]. The rendered human images in the target space provide explicit conditional information, reducing the difficulty of transforming information from the reference image space to the target space. In addition, we encode camera pose information into a Plücker ray representation [113], serving as a conditional signal of relative camera poses. To further enhance the spatial and temporal consistency, we incorporate a class embedding to utilize attention modules for novel view and pose synthesis tasks. For the second challenge, we observe a discrepancy between the latent and image space, attributed to the Variational Autoencoder (VAE) [72] used in diffusion models. Inspired by the Radiance field rendering loss employed in 3D object diffusion models [88], we construct an image-level loss by mapping the noise latent to the image space through a VAE decoder. This ensures consistent optimizations in both latent and image spaces. With the incorporated modules, our HumanGif successfully recovers fine-grained information from the reference image and achieves the best perceptually realistic, view-consistent, and temporally coherent results. Our main contributions are as follows:
@@
    To learn perpetually realistic, view-consistent, and temporally aligned 3D human avatars, we render human images in the target space, along with Plücker ray representation to serve as a more spatially and temporally aligned conditioning signal.
@@
    Extensive experiments on human datasets demonstrate that HumanGif outperforms baseline methods in perceptual quality of novel view and pose synthesis.
@@ 2 Related Work
Novel View/Pose Synthesis from a Single Human Image. Although 2D human modelling [2, 75, 109, 86, 31, 64, 65, 32] has made substantial progress, it is still challenging to synthesizing 3D humans from monocular inputs, especially for a single human image input [105, 106, 40, 79, 30, 77, 10, 133, 7, 8, 56, 41, 147, 130, 129, 5, 25]. To complement the information missing from a single image input, these approaches [126, 55, 91, 122, 20, 23, 144, 44, 29, 15, 14] typically leverage text-to-image (T2I) diffusion models, such as Stable Diffusion [101], by employing subject-specific Score Distillation Sampling (SDS) [96] or fine-tuning a T2I diffusion model conditioned on a front-view image using multi-view human datasets or 3D scan datasets. This enables the synthesis of human images from unseen viewpoints, such as generating a back view from a given front view. To further learn view-consistent and temporally aligned human avatars from a single image, recent research [51, 110] leverages the strong modeling capabilities of deep neural networks to learn generalizable features for Neural Radiance Field (NeRF) or diffusion models, enabling the synthesis of novel views and poses in a feed-forward manner. However, the scale of multi-view human datasets limits their generalization ability. In this work, we propose to utilize the generative prior from diffusion models to produce perpetually realistic, view-consistent, and temporally coherent human avatars from a single image.
@@
Diffusion Models for 2D Character Animation. 2D Character Animation aims at generating temporally coherent animation videos from one or more static human images [3, 12, 16, 31, 62, 97, 100, 108, 111, 112, 136, 138, 141, 146]. With advancements in text-to-image diffusion models, recent studies [9, 67, 121, 18, 132, 49, 149, 143, 69, 81] have investigated the use of diffusion models for 2D character animation. These approaches normally use a reference human image and a sequence of target pose images (e.g., OpenPose [13], DWPose [135], DensePose [35], or SMPL pose [82]) as conditional inputs to pose encoders or ControlNet [140], generating a sequence of target images aligned with the reference human image and target pose images. Motivated by the success of diffusion-based methods in 2D character animation, we adapt these methodologies for 3D/4D human modeling from a single image. Human4DiT [110] introduced a diffusion model that generates multi-view human animation videos from a single image, utilizing factorized image, view, and temporal modules. However, the code for this model is not publicly available. We identified that our baseline multi-view video diffusion model could not capture fine-grained details and maintain multi-view consistency. We show that integrating a generative prior and our proposed modules into the diffusion pipeline effectively resolves these challenges. A concurrent work GAS [83] also utilizes diffusion models to enhance the novel view and pose synthesis performance from a Human NeRF module.
@@ 3 Our Approach
Our proposed HumanGif, as illustrated in Fig. 2, learns a single-view conditioned human diffusion model for novel view and pose synthesis by utilizing a generative prior from Stable Diffusion. Specifically, given a reference human image and a target human pose sequence (estimated from a human video or generated from other modalities like text, or audio), HumanGif aims to predict a sequence of human pose images aligned with the person in the reference image and motions observed in the target human pose sequence.
Figure 2: HumanGif Framework. Given a single input human image and a target pose sequence, our HumanGif produces a sequence of target images aligned with the input image and target human poses. To synthesize view-consistent and temporally coherent outputs, our HumanGif proposes incorporating a generative prior, a Human NeRF module, Plücker ray representation, image-level loss, view/temporal attention.
@@ 3.1 Preliminary
3.2 Single-view Human Diffusion Model
@@
Pose Encoder. There are various options for defining target poses, including OpenPose, DensePose, DWPose, and SMPL parametric poses. We follow [149] to estimate SMPL mesh for the reference image, animate the SMPL mesh with the target SMPL pose parameters, and render normal images as the target guidance signal. The SMPL normal pose images are encoded through a Pose Encoder [49], which contains four convolution layers. The Pose Encoder output is added to the noise latent to provide view and pose guidance.
Table 1: Quantitative comparison of our HumanGif and baseline methods on the RendePeople, DNA-Rendering, THuman2.1 and TikTok datasets. * denotes that checkpoints released from the original work are used for performance evaluation.
@@
Method THuman2.1 & TikTok
Novel View Novel Pose
[…]
Champ [149] 3.36E-04 6.858 0.660 0.444 1.23E-04 13.395 0.727 0.275 35.31
SHERF [51] 1.07E-05 25.148 0.935 0.071 7.63E-05 16.930 0.715 0.257 20.46
Animate Anyone [49] 1.61E-05 21.487 0.932 0.054 9.25E-05 14.207 0.760 0.238 42.89
Champ [149] 1.13E-05 23.828 0.945 0.040 8.60E-05 14.761 0.768 0.231 32.93
HumanGif (Ours) 9.71E-06 25.966 0.956 0.029 8.77E-05 15.044 0.772 0.223 18.39
@@
Human NeRF. We observe two challenges by adopting human diffusion models from 2D character animation tasks: (a) difficulty in capturing fine-grained details (e.g., logos on T-shirts), even when they are present in the reference image, and (b) limited capability to maintain consistency across multiple views and poses. One underlying reason is that the reference image is encoded into a latent representation and processed by ReferenceNet to extract features, resulting in information loss during this process. Inspired by prior research in 3D human reconstruction methods [94, 51, 19, 92, 131, 90, 124, 115, 61, 63, 120, 74, 33, 23, 145, 54], we introduce a more spatially and temporally aligned conditioning signal by rendering human images in the target space.
@@
The input to the Human NeRF module is a single human image $\mathbf{I}^{\text{ref}}$ along with its corresponding camera parameters $\bm{P}^{\text{ref}}$ and SMPL pose parameter $\bm{\theta}^{\text{ref}}$ and shape parameter $\bm{\beta}^{\text{ref}}$. The module outputs the rendered human feature image in the target camera view $\bm{P}^{\text{tgt}}$, corresponding to the target SMPL pose $\bm{\theta}^{tgt}$ and shape $\bm{\beta}^{\text{tgt}}$. Specifically, in the target space, we cast rays passing through the camera origin and image pixels, and sample points $\bm{x}^{\text{tgt}}$ along the cast rays. These points $\bm{x}^{\text{tgt}}$ are transformed into the canonical space $\bm{x}^{c}$ using inverse Linear Blend Skinning (LBS). Subsequently, hierarchical 3D-aware features are queried from their respective feature extraction modules. The queried features are concatenated and passed to the NeRF decoder to predict the density $\sigma$ and feature $\bm{c}$ for each sampled point. The final pixel features are rendered in the target space through volume rendering, integrating the density and feature values of the sampled 3D points along the rays in the target space. The details of extracting 3D-aware features are described in the appendix.
Plücker Ray Representation. Camera pose information is important for novel view and pose synthesis, as relative camera poses are attributed to the reference-to-target and target-to-target transformation. We encode camera pose information into a Plücker ray representation [113] and add it to the input noise signal, serving as a conditional signal of camera poses.
@@
View/Temporal Module. To learn view-consistent and temporally coherent 3D human avatars, we further integrate a view/temporal attention layer after the spatial-attention and cross-attention components within the Res-Trans block of Denosing UNet. Instead of using two separate attention modules for synthesizing views and poses, we adopt a unified attention module but different class embeddings for the two tasks. Inspired by the efficient temporal attention layer adopted in [49, 36], we utilize the same architecture for our view/temporal attention layer. Furthermore, inspired by [42], we extend the original self-attention in the Denoising UNet to be 3D attention. The 3D attention learns view/temporal information by conducting 3D attention in a selected number of feature maps from nearby views or time steps.